Re: Shift report 20000714 00:00-08:00

From: Konstantin Olchanski (olchansk@ux1.phy.bnl.gov)
Date: Fri Jul 14 2000 - 13:36:26 EDT

  • Next message: http account: "Shift report 20000714 08:00-16:00"

    On Fri, Jul 14, 2000 at 07:48:14AM -0400, http account wrote:
    > 
    > ------------------------------------------------------------
    > Date: Fri Jul 14 07:48:14 2000, Shift: 00:00-08:00
    > Supervisor: Michael Murray, with: all alone with only the deer to talk to
    > Log book#2 pages 151 - 155
    > 
    > Shift Summary:
    > ------------------------------------------------------------
    > ...
    > TOFW lost the TDC in slot 10 and this confused the DAQ
    > so we got no TOFW data through the night. ...
    
    
    The TDC developed a bad bit (0x4000) and has been fixed by reseating
    the TDC in the crate (take it out, put it back, do not forget to
    turn off the power). The unpacking software will be modified to treat
    this type of error as non-fatal, so the data can still be looked at when
    one of the TDCs goes bad.
    
    
    -- 
    Konstantin Olchanski
    Physics Department, Brookhaven National Laboratory, Long Island, New York
    olchansk@bnl.gov
    



    This archive was generated by hypermail 2b29 : Fri Jul 14 2000 - 13:40:07 EDT